# Breakdown of broken-symmetry approach to exchange interaction

### v1

@@ III.3.1 MC DFT approaches
where the first term is the CAS energy (20) in its conventional form (without rescaling), while the second term is the correlation energy in which the conventional spin densities are replaced by combinations of total CAS density $\rho(\mathbf{r})$ and on-top pair density $P(\mathbf{r},\mathbf{r})$ [71]. As appropriate $E^{\text{C}}$, a Colle-Salvetti [72, 73] or Lee-Yang-Parr [74] correlation functional should be used. In order to avoid a double counting of dynamical correlation energy covered by the first term of (21), $E^{\text{CAS-DFT}}_{\text{C}}[\rho,P]$ is evaluated with local rescaling factors [70]. This rescaling can be neglected when a few configurations are mixed in the CASSCF wave function, which is certainly the case here. Indeed, given the relative weakness of the exchange coupling in most magnetic complexes, we will only need to consider singly and doubly excited configurations from reference one(s) resulting in their limited amount even for a large space of active magnetic and ligand orbitals.
@@
A related version of CAS-DFT is the actively developed multiconfigurational pair-density functional theory (MC-PDFT) [75, 76]. It is currently implemented in OpenMolcas [77] with a plethora of on-top functionals, corresponding to translated exchange-correlation functionals, and different versions of MC calculations. It has been successfully applied to the calculation of relative energy levels in exchange-coupled systems [78], the calculation of singlet-triplet splittings in main-group and organic systems [79, 80, 81] and the relative spin-state energetics of coordinated metal ions [82, 83].
@@
Contrary to these methods which do not involve the KS density, the REKS functional is a linear combination KS energies corresponding to different electronic configurations of active electrons. The coefficients of this combination are expressed via the fractional occupation numbers in the total REKS density trough model considerations [65, 66, 67]. The weak point of this approach is that it is designed for very small active spaces and cannot be easily extended over, e.g., CAS(4,4). Note that the latter will be already sufficient for the calculation of exchange parameters in symmetric dimers with one unpaired electron per site, for which the expressions derived in Sec. III.2 and Appendix C can be applied directly. The densities corresponding to different electronic configurations of active electrons are calculated as described in Appendix C.2. As an example, Eq. (48) gives the total density for the ground configuration $\Phi_{1}$. One should have in mind that the self-consistent procedure involves variation of orbital coefficients defining the active orbitals only. The same refers also to other approaches mentioned above.
@@
One should note that the MC DFT approaches have been straightforwardly applied to the calculation of exchange parameters in organic materials [84, 61] and complexes (an overview of earlier work can be found in Ref. 85). They generally produced results comparable in accuracy with BS DFT calculations. Thus CASSCF(2,2) calculations of magnetic coupling in Cu(II) binuclear complexes by the REKS method have shown that imposing strict spin symmetry does not improve the BS DFT evaluation of exchange parameters [86]. While the active space in these calculations was restricted to magnetic orbitals only, we stress that including specially designed active ligand orbitals are expected to improve the predictibility of $J$ in such calculations, as they would certainly improve the BS DFT results according to the present study.
IV Conclusion
@@
In this work, we prove the breakdown of the broken-symmetry approach for the evaluation of exchange parameters by applying it to a generic three-site model. We show that this breakdown originates from an artificial constraint on multiconfigurational state imposed by the broken-symmetry determinant. The error becomes especially pronounced in the case of strong covalency between magnetic centers and the bridging ligand. To cure this drawback, we propose a calculational scheme based on a minimal multiconfigurational extension of the BS approach. An example of such an economical employment of the CI space for the description of realistic systems is the recently developed GS-ROCIS method [87].
@@
As active orbitals in the MC calculations, the proposed method employs effective magnetic and bridging ligand type orbitals, whose construction and self-consistent determination is outlined in detail. This approach can be used with any a variety of quantum chemistry software involving MC and BS calculations, in particular, with any version of existent MC DFT code, the only required modification being the implementation of optimization of the coefficients defining the active orbitals. Besides possible improvement in the prediction of exchange parameters, we expect this approach to help resolving the issue related to the strong variation of the performance of a given exchange-correlation functional for evaluation of $J$ in different magnetic systems [21, 22], which variability seems to be less pronounced for other molecular properties calculated with DFT.
@@ Acknowledgement
N.I. was supported by Grant-in-Aid for Scientific Research (Grant No. 22K03507) from the Japan Society for the Promotion of Science, and Chiba University Open Recruitment for International Exchange Program. Z.H. was supported by the China Scholarship Council, and the financial support of research projects A-8000709-00-00, A-8000017-00-00, and A-8001894-00-00 of the National University of Singapore. A.M. A. M. acknowledges funding provided by the Research Council of Finland (grant no. 362649).
@@ Appendix A Solutions for the generic three-site model
The full CI and broken-symmetry HF treatment of the Hubbard model (1) is shown below.
@@ A.1.1 Ferromagnetic state
The basis for the ferromagnetic (spin triplet, $S=1$) states $|\text{F},M_{S};n,p\rangle$ are
@@
respectively. Here, $|ij\rangle$ and $|i\bar{j}\rangle$ etc. indicate Slater determinants, spin up and down are specified by without and with bar, “F” stands for ferromagnetic state, $M_{S}$ the $z$ component of total spin, $n$ distinguishes the $S,M_{S}$ states, $p$ (=±absentplus-or-minus=$\pm$= ±) the parity of the spatial part (symmetric or antisymmetric). The Hamiltonian matrix is written as
@@
The order of the basis is $|0-\rangle$, $|1-\rangle$, $|1+\rangle$ (“F” and $M_{S}$ are omitted). The ground energy is obtained from the $2\times 2$ block, and the ground state is expressed as
@@ A.1.2 Antiferromagnetic state
The symmetrized antiferromagnetic (singlet) states $|\text{AF};n,p\rangle$ are
@@
in the order of the basis $|0+\rangle$, $|1+\rangle$, $|2+\rangle$, $|3+\rangle$, $|1-\rangle$, $|2-\rangle$ (“AF” is omitted). The ground energy is obtained from the spatially symmetric part (the $4\times 4$ block), and the ground antiferromagnetic state is written as
@@
In Fig. 2 in the main text, the weight of each configuration $w_{i}=C_{i}^{2}$ in function of $\Delta/t$ is displayed.
@@ A.2.1 Ferromagnetic state
Since both the high-spin full CI and HF wave functions contain one variational parameter, the exact high-spin state can be obtained by HF approach.
Figure 5: $\langle S^{2}\rangle_{\text{UHF}}$ with respect to $t^{\prime}/U$. The same parameters as Fig. 3(a) are used.
A.2.2 Unrestricted low-spin state
The molecular orbitals used for the unrestricted HF (UHF) method are written as
@@ A.2.2 Unrestricted low-spin state
where the coefficients $C_{1},C_{2},C_{l}$ are real and the molecular orbitals are normalized. The low-spin UHF wave function,
$|\Psi_{\text{UHF}}\rangle$ $=$ $|\psi^{\uparrow}\bar{\psi}^{\downarrow}\rangle$
@@
This expression shows both the ferro- and antiferromagnetic configurations are included in $|\Psi_{\text{UHF}}\rangle$. Note that the exact low-spin states are expressed by three parameters, while there are only two parameters in the UHF states. Based on the wave function (LABEL:Eq:WF_UHF_config), the UHF energy and $\langle\hat{\bm{S}}^{2}\rangle_{\text{UHF}}$ are calculated as, respectively,
$E_{\text{UHF}}$ $=$ $4tC_{l}(C_{1}+C_{2})+4t^{\prime}C_{1}C_{2}+2\Delta C_{l}^{2}$ (33)
@@
$\langle\hat{\bm{S}}^{2}\rangle_{\text{UHF}}$ $=$ $\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}\left[\frac{(C_{1}^{2}-C_{2}^{2})^{2}}{2}+(C_{1}-C_{2})^{2}C_{l}^{2}\right].\qquad$ (34)
Calculated $\langle\hat{\bm{S}}^{2}\rangle_{\text{UHF}}$ is shown in Fig. 5.
@@ A.2.3 Spin symmetry adapted state
We define the AF part of the AF part of UHF wave function as the spin symmetry adapted (SA) HF wave function:
$|\Psi_{\text{SAHF}}\rangle$ $=$ $\frac{1}{\sqrt{2(1+S_{M}^{2})}}\left(|\psi^{\uparrow}\bar{\psi}^{\downarrow}\rangle+|\psi^{\downarrow}\bar{\psi}^{\uparrow}\rangle\right),$ (35)
where $S_{M}=2C_{1}C_{2}+C_{l}^{2}$ is the overlap between $\psi^{\uparrow}$ and $\psi^{\downarrow}$ (31). The energy for the SA-HF state is
$E_{\text{SAHF}}$ $=$ $\frac{1}{1+S_{M}^{2}}\Big{[}4t(1+S_{M})(C_{1}+C_{2})C_{l}$ (36)
@@
$+2\Delta C_{l}^{2}(1+S_{M})+4UC_{1}^{2}C_{2}^{2}+2U^{\prime}C_{l}^{4}\Big{]}.\qquad$
$J$ was calculated using a modified Yamaguchi’s formula in which $E_{\text{UHF}}$ and $\langle S^{2}\rangle_{\text{UHF}}$ are replaced by $E_{\text{SA}}$ and $\langle S^{2}\rangle_{\text{SA}}=0$, respectively.
@@ Appendix B Expressions for the magnetic force theorem
Here, we show the detailed expressions used for the magnetic force theorem calculations of $J$. The rotations of the quantum spins of the broken symmetry state results in
@@
In terms of the configurations, we have
@@
The energy expectation value for $\hat{R}(\theta)|\psi^{\uparrow}\bar{\psi}^{\downarrow}\rangle$ is
$E(\theta)$ $\approx E_{\text{UHF}}+\theta^{2}\left[-4(C_{1}+C_{2})C_{l}(2C_{1}C_{2}+C_{l}^{2})t\right.$
@@ C.2 MC calculation in the electron-hole representation
For the restricted doubly occupied orbitals $\psi_{\mu}$ we pass from electron to the hole representation. In the language of second quantization [50] the electron creation is replaced by hole annihilation and vice versa:
@@ C.2.2 The Hamiltonian in the electron-hole representation
where $\hat{H}_{a}$ is the electronic part given by Eq. (49) in which the summations over orbital indices exclude the orbitals $\psi_{\mu}$, and $\hat{H}_{b}$ is the Hamiltonian for holes (in the following formulas the hole orbitals are denoted by Greek letters):
@@
The above expressions were derived for real orbitals implying usual symmetry relations for matrix elements (indices refer to all orbitals:

### v2

@@ V.3.1 MC DFT approaches
where the first term is the CAS energy (25) in its conventional form (without rescaling), while the second term is the correlation energy in which the conventional spin densities are replaced by combinations of total CAS density $\rho(\mathbf{r})$ and on-top pair density $P(\mathbf{r},\mathbf{r})$ [82]. As appropriate $E^{\text{C}}$, a Colle-Salvetti [83, 84] or Lee-Yang-Parr [85] correlation functional should be used. In order to avoid double counting of dynamical correlation energy covered by the first term of (26), $E^{\text{CAS-DFT}}_{\text{C}}[\rho,P]$ is evaluated with local rescaling factors [81]. This rescaling can be neglected when a few configurations are mixed in the CASSCF wave function, which is certainly the case here. Indeed, given the relative weakness of the exchange coupling in most magnetic complexes, we will only need to consider singly and doubly excited configurations from reference one(s), resulting in their limited amount even for a large space of active magnetic and ligand orbitals.
@@
A related version of CAS-DFT is the actively developed multiconfigurational pair-density functional theory (MC-PDFT) [86, 87]. It is currently implemented in OpenMolcas [88] with a plethora of on-top functionals corresponding to translated exchange-correlation functionals and different versions of MC calculations. It has been successfully applied to the calculation of relative energy levels in exchange-coupled systems [89], the calculation of singlet-triplet splittings in main-group and organic systems [90, 91, 92] and the relative spin-state energetics of coordinated metal ions [93, 94].
@@
Contrary to these methods, which do not involve the KS density, the REKS functional is a linear combination of KS energies corresponding to different electronic configurations of active electrons. The coefficients of this combination are expressed via the fractional occupation numbers in the total REKS density through model considerations [76, 77, 78]. The weak point of this approach is that it is designed for tiny active spaces and cannot be easily extended over, e.g., CAS(4,4). Note that the latter will be already sufficient for the calculation of exchange parameters in symmetric dimers with one unpaired electron per site, for which the expressions derived in Sec. V.2 and Appendix C can be applied directly. The densities corresponding to different electronic configurations of active electrons are calculated as described in Appendix C.2. As an example, Eq. (55) gives the total density for the ground configuration $\Phi_{1}$. One should keep in mind that the self-consistent procedure involves a variation of orbital coefficients defining the active orbitals only. The same refers also to other approaches mentioned above.
@@
One should note that the MC DFT approaches have been straightforwardly applied to the calculation of exchange parameters in organic materials [95, 72] and complexes (an overview of earlier work can be found in Ref. 96). They generally produced results comparable in accuracy with BS DFT calculations. Thus, CASSCF(2,2) calculations of magnetic coupling in Cu(II) binuclear complexes by the REKS method have shown that imposing strict spin symmetry does not improve the BS DFT evaluation of exchange parameters [97]. While the active space in these calculations was restricted to magnetic orbitals only, we stress that including specially designed active ligand orbitals is expected to improve the predictability of $J$ in such calculations, as they would certainly improve the BS DFT results according to the present study.
VI Conclusion
@@
In this work, we prove the breakdown of the broken-symmetry approach for the evaluation of exchange parameters by applying it to a generic three-site model. We show that this breakdown originates from an artificial constraint on the multiconfigurational state imposed by the broken-symmetry determinant. It is also found that other single-reference based approaches do not cure this drawback. The error becomes significantly pronounced in the case of strong covalency between magnetic centers and the bridging ligand. To cure this drawback, we propose a calculational scheme based on a minimal multiconfigurational extension of the BS approach. An example of such an economical employment of the CI space for the description of realistic systems is the recently developed GS-ROCIS method [98].
@@
As active orbitals in the MC calculations, the proposed method employs effective magnetic and bridging ligand-type orbitals, whose construction and self-consistent determination are outlined in detail. This approach can be used with a variety of quantum chemistry software involving MC and BS calculations, in particular, with any version of the existing MC DFT code, the only required modification being the implementation of optimization of the coefficients defining the active orbitals. Besides possible improvement in the prediction of exchange parameters, we expect this approach to help resolve the issue related to the strong variation of the performance of a given exchange-correlation functional for evaluation of $J$ in different magnetic systems [21, 22], which variability seems to be less pronounced for other molecular properties calculated with DFT.
@@ Acknowledgement
N.I. was supported by Grant-in-Aid for Scientific Research (Grant No. 22K03507) from the Japan Society for the Promotion of Science, and Chiba University Open Recruitment for International Exchange Program. Z.H. was supported by the China Scholarship Council, and the financial support of research projects A-8000709-00-00, A-8000017-00-00, and A-8001894-00-00 of the National University of Singapore. A.M. acknowledges funding provided by the Research Council of Finland (grant no. 362649).
@@ Appendix A Solutions for the generic three-site model
The full CI and HF treatments of the three-center Hubbard model (1) are shown below.
@@ A.1.1 Ferromagnetic state
The basis for the ferromagnetic (spin-triplet, $S=1$) states $|\text{F},M_{S};n,p\rangle$ are
@@
respectively. Here, $|ij\rangle$ and $|i\bar{j}\rangle$ etc. indicate Slater determinants, spin up and down are specified by without and with bar, “F” stands for ferromagnetic state, $M_{S}$ the $z$ component of the total spin, $n$ distinguishes the states characterized by the same $S$ and $M_{S}$, $p$ (=±absentplus-or-minus=$\pm$= ±) the parity of the spatial part (symmetric or antisymmetric).
The Hamiltonian matrix is written as
@@
with the ferromagnetic basis in the order of $|0-\rangle$, $|1-\rangle$, $|1+\rangle$ (“F” and $M_{S}$ are omitted). The ground energy is obtained from the $2\times 2$ block of Eq. (29), and the ground state is expressed as
@@ A.1.2 Antiferromagnetic state
The symmetrized antiferromagnetic (spin-singlet, $S=0$) states $|\text{AF};n,p\rangle$ are
@@
in the order of the basis $|0+\rangle$, $|1+\rangle$, $|2+\rangle$, $|3+\rangle$, $|1-\rangle$, $|2-\rangle$ (“AF” is omitted). The ground energy is obtained from the spatially symmetric part (the $4\times 4$ block). The ground antiferromagnetic state is written as
@@
The exact antiferromagnetic ground state of the present model (1) has three parameters considering the normalization condition.
@@ A.2.1 Ferromagnetic state
Since both the high-spin full CI and HF wave functions contain one variational parameter, the HF wave function can describe the exact high-spin state.
A.2.2 Broken symmetry low-spin state
The molecular orbitals used for the BS-HF (or unrestricted HF, UHF) method are written as
@@ A.2.2 Broken symmetry low-spin state
where the coefficients $C_{1},C_{2},C_{l}$ are real, and the molecular orbitals are normalized. The BS-HF wave function is
$|\Psi_{\text{BS}}\rangle$ $=$ $|\psi^{\uparrow}\bar{\psi}^{\downarrow}\rangle$
@@
This expression shows that both ferro- and antiferromagnetic configurations are included in $|\Psi_{\text{BS}}\rangle$. Note that the exact low-spin states have three variational parameters, while the BS states contain only two parameters.
Based on the wave function (LABEL:Eq:WF_UHF_config), the BS-HF energy and $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}$ are calculated as, respectively,
$E_{\text{BS}}$ $=$ $4tC_{l}(C_{1}+C_{2})+4t^{\prime}C_{1}C_{2}+2\Delta C_{l}^{2}$ (39)
@@
$\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}$ $=$ $\langle\hat{\bm{S}}^{2}\rangle_{\text{F}}\left[\frac{(C_{1}^{2}-C_{2}^{2})^{2}}{2}+(C_{1}-C_{2})^{2}C_{l}^{2}\right].\qquad$ (40)
We determine the molecular orbital coefficients by numerically minimizing $E_{\text{BS}}$. The calculated $\langle\hat{\bm{S}}^{2}\rangle_{\text{BS}}$ is shown in Fig. 3.
@@ A.2.3 Spin symmetry adapted state
The spin symmetry adapted (SA) HF wave function is the AF ($S=0$) part of the BS wave function (LABEL:Eq:WF_UHF_config):
$|\Psi_{\text{SA}}\rangle$ $=$ $\frac{1}{\sqrt{2(1+S_{M}^{2})}}\left(|\psi^{\uparrow}\bar{\psi}^{\downarrow}\rangle+|\psi^{\downarrow}\bar{\psi}^{\uparrow}\rangle\right)$
$=$ $\frac{1}{\sqrt{1+S_{M}^{2}}}\Big{[}(C_{1}^{2}+C_{2}^{2})|\text{AF};0+\rangle$
$+$ $\sqrt{2}(C_{1}+C_{2})C_{l}|\text{AF};1+\rangle+2C_{1}C_{2}|\text{AF};2+\rangle$
$+$ $\sqrt{2}C_{l}^{2}|\text{AF};3+\rangle\Big{]},$ (42)
where $S_{M}=\langle\psi^{\uparrow}|\psi^{\downarrow}\rangle=2C_{1}C_{2}+C_{l}^{2}$. Since the right-hand side contains only the AF configurations, the spin expectation value for $|\Psi_{\text{SA}}\rangle$ is 0.
The energy for the SA-HF state is
$E_{\text{SA}}$ $=$ $\frac{1}{1+S_{M}^{2}}\Big{[}4t(1+S_{M})(C_{1}+C_{2})C_{l}$ (43)
@@
$+2\Delta C_{l}^{2}(1+S_{M})+4UC_{1}^{2}C_{2}^{2}+2U^{\prime}C_{l}^{4}\Big{]}.$
We obtain the molecular orbital coefficients by minimizing $E_{\text{SA}}$.
@@ Appendix B Expressions for the magnetic force theorem
Here, we show the detailed expressions used for the magnetic force theorem calculations of $J_{\text{MFT}}$. The rotations of the quantum spins of the broken symmetry state (A.2.2) results in
@@
Here, $\hat{R}(\theta)$ corresponds to Eq. (9). In terms of the configurations, the rotated broken symmetry state is
@@
The energy expectation value for $\hat{R}(\theta)|\psi^{\uparrow}\bar{\psi}^{\downarrow}\rangle$ is, for small $\theta$,
$E(\theta)$ $\approx E_{\text{BS}}+\theta^{2}\left[-4(C_{1}+C_{2})C_{l}(2C_{1}C_{2}+C_{l}^{2})t\right.$
@@ C.2 MC calculation in the electron-hole representation
For the restricted doubly occupied orbitals $\psi_{\mu}$, we pass from the electron to the hole representation. In the language of second quantization [61] the electron creation is replaced by hole annihilation and vice versa:
@@ C.2.2 The Hamiltonian in the electron-hole representation
where $\hat{H}_{a}$ is the electronic part given by Eq. (56) in which the summations over orbital indices exclude the orbitals $\psi_{\mu}$, and $\hat{H}_{b}$ is the Hamiltonian for holes (in the following formulas, the hole orbitals are denoted by Greek letters):
@@
The above expressions were derived for real orbitals, implying usual symmetry relations for matrix elements (indices refer to all orbitals: